OpenStudy (anonymous):

a tugboat goes 100 miles upstream in 10 hours. The return trip takes 5 hours. Find the speed of the tugboat without a current and the speed of the current. Speed is what and speed of current is

OpenStudy (ganpat):

speed upstream = 100/10 = 10 miles/hr Speed downstream = 100/5 = 20 miles/hr so speed of boat = (20 + 10 )/ 2 = 15 miles/ hr and speed of current = (20 - 10) / 2 = 5 miles/ hr..

OpenStudy (anonymous):

thanks

OpenStudy (ganpat):

understood ??

OpenStudy (anonymous):

yes speed is 15 and speed of current is 5 miles

OpenStudy (ganpat):

h ah ah, not the answer, the formula stuff ??

OpenStudy (anonymous):

what do youi mean do I have to write the whole formula

OpenStudy (anonymous):

General solution: \[2 \frac {d}{u+v} = 2(5) = 10 = \frac {d}{u-v}\]

OpenStudy (anonymous):

u = 3v

OpenStudy (anonymous):

Sorry @Ganpat thought you were done.

OpenStudy (ganpat):

i mean, y i have written something like that, from the whole book of formulas..

OpenStudy (ganpat):

@telliott99 : No issues brother !! every approach should be mentioned, so that the user may use the optimized one :) :)

OpenStudy (anonymous):

I am so confusd

OpenStudy (anonymous):

Sorry, just go with his answer above. Come back when you need a general equation.

OpenStudy (anonymous):

Ganpat do I put the whole equation in

OpenStudy (ganpat):

over here ?? not required, but try to be complete in notes.. so that next time, you might understand, how to arrive to a solution.. I just wanted to make sure, u understood this..

OpenStudy (anonymous):

so what d I get for t speed and the speed of current

OpenStudy (anonymous):

See the very first post in this thread.

OpenStudy (anonymous):

I did i it 15mph for speed and 5 mlesfor the current

OpenStudy (ganpat):

u r right.. :)

OpenStudy (anonymous):

thanks